OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

y=3x+b -2=3*3+b so b=-11 y=3x-11